OpenStudy (candycanesschick):

Can someone help me with these 5 math questions. I will give you a medal.

OpenStudy (pianogirl3):

What are the questions?

OpenStudy (candycanesschick):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal?

OpenStudy (candycanesschick):

Data set Mean Mean absolute deviation (MAD) 1 10.7 2.4 2 13.7 2.5

OpenStudy (candycanesschick):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {43, 42, 56, 62, 72} Data Set 2: {76, 57, 81, 51, 70}

OpenStudy (pianogirl3):

First you need to find the mean, do you know how to do that?

OpenStudy (candycanesschick):

Is that adding all of the numbers then dividing by how many it is

OpenStudy (pianogirl3):

yes, so do that first

OpenStudy (candycanesschick):

For both sets ?

OpenStudy (blm1):

yes

OpenStudy (pianogirl3):

That's what it looks like

OpenStudy (candycanesschick):

217.4 first set and second set 279

OpenStudy (pianogirl3):

43 +42 + 56 + 62 + 72 = 275 275 / 5 =55

OpenStudy (candycanesschick):

Now divide 275 and 55?

OpenStudy (blm1):

No, you found the mean wrong, so she was showing you how to properly find it :)

OpenStudy (pianogirl3):

Now you have to find the distance between each number and the mean. Like 76 -55 = 21, do this for all of the numbers.

OpenStudy (pianogirl3):

Once you find the each number subtracted by the mean you add all of them together. Then divide again by 5.

OpenStudy (candycanesschick):

Is the answer 1.2?

OpenStudy (pianogirl3):

I completed the steps by subtracting 55 from each number. So I got -12, -13, 1, 7, 17. Add together = 0. 0/ 5 (the amount of numbers) = 0

OpenStudy (pianogirl3):

Does this help?

OpenStudy (candycanesschick):

Are you with k12? This is one of their lessons. And this is what the lesson looks like.

OpenStudy (pianogirl3):

No

OpenStudy (candycanesschick):

wrong link

OpenStudy (candycanesschick):

It wont let me post it

OpenStudy (pianogirl3):

Ok, did I help you?

OpenStudy (candycanesschick):

Yes, I've kind of got how to do it

OpenStudy (pianogirl3):

I would love for you to fan and medal me. I have to go, bye :)